hi please can you solve: 4logx-6=0 and show working with notes if possible please and thank you
Answer by mananth(16946)

4logx -6=0
4logx=6
logx=6/4
logx=3/2
x= 10^(3/2)
x= sqrt(10^3)
x= sqrt(1000)
x=10 sqrt(10)